SITUATIONAL ANALYSIS:


Two fishers went spearfishing, one using gas cylinder and the other without using any breathing apparatus. At depths of 100ft, fishers are subject to approximately 5atm. As they go deeper, the fisher without breathing apparatus is experiencing difficulty breathing while the other feels normal. Assuming a typical gas cylinder contains 50 grams of O2 and 250 grams of He and has a volume of 8.0 L. What is the partial pressure of each gas at 25 OC? What is the total pressure inside the cylinder? Why does the fisher without breathing apparatus experience difficulty breathing as he went deeper?


1
Expert's answer
2021-11-30T06:04:01-0500

Moles of "O_2=\frac{50}{32}=" 1.5625mole


Moles of "He" "=\frac{250}{4}=62.5moles"


Molarity of He "= \frac{62.5}{8}=7.8125M"



In freshwater, 34 ft. (10.4 m) equals 14.7 psi or 0.432 psi per foot of freshwa- ter (ffw). Thereafter, for every 34 ft. of additional depth in freshwater, the hydrostatic pressure increases by one atmosphere